OpenStudy (lgbasallote):

find f \[f''(x) = x^{-2}, \quad x > 0,\quad f(1) = 0, \quad f(2) = 0\]

myininaya (myininaya):

To find f you need to first find f' the find f' given f'' integrate both sides

OpenStudy (lgbasallote):

hmm.. \[f'(x) = \frac 1x + c\] ??

myininaya (myininaya):

almost

myininaya (myininaya):

you are missing a certain constant multiple

OpenStudy (lgbasallote):

ahh \[f'(x) = -\frac 1x + c\]

myininaya (myininaya):

yes :) now to find the constant hmmm....you are missing a certain initial condition to do that .... your question doesn't make sense .... you need one of those to be f'(something)=another something

OpenStudy (lgbasallote):

or maybe i should do it \[f'(x) = -\frac 1x + c_1\]

OpenStudy (lgbasallote):

i might have typoed

OpenStudy (zarkon):

integrate again

OpenStudy (lgbasallote):

ah yes i did. it's f'(2) not just f(2)

OpenStudy (zarkon):

get a system of 2 eau and 2 unknowns

OpenStudy (lgbasallote):

hmm \[f(x) = -\ln x + c_1 x + c_2\] yes?

myininaya (myininaya):

ok great. you can find that constant by using f(1)=0 and then do what zarkon says to find f

OpenStudy (zarkon):

you can do the problem with two given values of f

myininaya (myininaya):

or you can wait to find the first constant whatever

OpenStudy (lgbasallote):

i suppose x > 0 is just there to note that -ln x exists?

myininaya (myininaya):

oh wait.... i guess you can do it with f(something1)=another something1 and f(something2)=another something2 oops

OpenStudy (lgbasallote):

f'(2) = 0 so f'(2) = -1/2 + c_1 = 0 does this mean c_1 is 1/2?

myininaya (myininaya):

yes adding 1/2 to both sides solves that equation for c_1

OpenStudy (lgbasallote):

then f(1) = -ln (1) + 1/2 x + c_2 = 0 so c_2 is -1/2?

myininaya (myininaya):

x is 1 so you have -ln(1)+1/2(1)+c_2=0 and yes

OpenStudy (lgbasallote):

oh. yeah...forgot to sub 1 into x the second time

OpenStudy (lgbasallote):

so \[f(x) = -\ln x + \frac 12 x - \frac 12\] ??

myininaya (myininaya):

tep

OpenStudy (lgbasallote):

nice. thanks

OpenStudy (lgbasallote):

i just noticed this was my 1000th question